# M-learner:A Flexible And Powerful Framework To Study Heterogeneous Treatment Effect In Mediation Model

### v1

@@ A.5.1 Without mediators
For each scenario , we estimate $\hat{g}_{0}(x)$ and $\hat{g}_{1}(x)$ using randomized clinical trial data. To facilitate visualization, we construct a synthetic grid of covariates where Cov1 and Cov2 vary from $-1.5$ to $1.5$ in increments of $0.05$, while all other covariates are drawn from a standard normal distribution. The estimated functions are then applied to this grid to compute the CATTE for each unit. We repeat this process across 100 simulated experiments and compute the average CATTE at each grid point. The resulting surface is interpolated using the R package "akima" and visualized to illustrate the spatial patterns in treatment heterogeneity. We visualize the results under a sample size of $1000$ for each of the four scenarios—Simple, Complex, Global, and Null—using Random Forest and XGBoost as base learners. The visualizations are presented in Appendix Figure A.10 - A.13.
@@
Figure A.10: Visualization of CATTE in the Simple scenario:each panel plots Cov1 (x-axis,$-1.4$ to $1.4$) and Cov2 (y-axis, $-1.4$ to $1.4$), with color representing the estimated CATTE magnitude.(a) visualization of RF learner result, (b) visualization of XGB learner result.
Figure A.11: Visualization of CATTE in the Complex scenario:each panel plots Cov1 (x-axis,$-1.4$ to $1.4$) and Cov2 (y-axis, $-1.4$ to $1.4$), with color representing the estimated CATTE magnitude.(a) visualization of RF learner result, (b) visualization of XGB learner result.
Figure A.12: Visualization of CATTE in the Global scenario:each panel plots Cov1 (x-axis,$-1.4$ to $1.4$) and Cov2 (y-axis, $-1.4$ to $1.4$), with color representing the estimated CATTE magnitude.(a) visualization of RF learner result, (b) visualization of XGB learner result.
Figure A.13: Visualization of CATTE in the Null scenario:each panel plots Cov1 (x-axis,$-1.4$ to $1.4$) and Cov2 (y-axis, $-1.4$ to $1.4$), with color representing the estimated CATTE magnitude.(a) visualization of RF learner result, (b) visualization of XGB learner result.
A.5.2 With a mediator
@@
For each scenario , we estimate $\hat{g}^{Y}_{1}$,$\hat{g}^{M}_{1}(x)$ and $\hat{g}^{M}_{0}(x)$ using randomized clinical trial data. To facilitate visualization, we construct a synthetic grid of covariates where Cov1 and Cov2 vary from $-1.5$ to $1.5$ in increments of $0.05$, while all other covariates are drawn from a standard normal distribution. The estimated functions are then applied to this grid to compute the CAITE for each unit. We repeat this process across 100 simulated experiments and compute the average CAITE at each grid point. The resulting surface is interpolated using the R package "akima" and visualized to illustrate the spatial patterns in treatment heterogeneity. We visualize the results under a sample size of $1000$ for each of the seven scenarios—Simple-All, Simple-Part, Complex-All, Complex-Part, Simple-Null1, Simple-Null2 and Simple-Global—using Random Forest and XGBoost as base learners. The visualizations are presented in Appendix Figure A.14 - A.20.
@@
Figure A.14: Visualization of CAITE in the Simple-All scenario:each panel plots Cov1 (x-axis,$-1.4$ to $1.4$) and Cov2 (y-axis, $-1.4$ to $1.4$), with color representing the estimated CAITE magnitude.(a) visualization of RF learner result, (b) visualization of XGB learner result.
Figure A.15: Visualization of CAITE in the Simple-Part scenario:each panel plots Cov1 (x-axis,$-1.4$ to $1.4$) and Cov2 (y-axis, $-1.4$ to $1.4$), with color representing the estimated CAITE magnitude.(a) visualization of RF learner result, (b) visualization of XGB learner result.
Figure A.16: Visualization of CAITE in the Complex-All scenario:each panel plots Cov1 (x-axis,$-1.4$ to $1.4$) and Cov2 (y-axis, $-1.4$ to $1.4$), with color representing the estimated CAITE magnitude.(a) visualization of RF learner result, (b) visualization of XGB learner result.
Figure A.17: Visualization of CAITE in the Complex-Part scenario:each panel plots Cov1 (x-axis,$-1.4$ to $1.4$) and Cov2 (y-axis, $-1.4$ to $1.4$), with color representing the estimated CAITE magnitude.(a) visualization of RF learner result, (b) visualization of XGB learner result.
Figure A.18: Visualization of CAITE in the Simple-Null1 scenario:each panel plots Cov1 (x-axis,$-1.4$ to $1.4$) and Cov2 (y-axis, $-1.4$ to $1.4$), with color representing the estimated CAITE magnitude.(a) visualization of RF learner result, (b) visualization of XGB learner result.
Figure A.19: Visualization of CAITE in the Simple-Null2 scenario:each panel plots Cov1 (x-axis,$-1.4$ to $1.4$) and Cov2 (y-axis, $-1.4$ to $1.4$), with color representing the estimated CAITE magnitude.(a) visualization of RF learner result, (b) visualization of XGB learner result.
Figure A.20: Visualization of CAITE in the Simple-Global scenario:each panel plots Cov1 (x-axis,$-1.4$ to $1.4$) and Cov2 (y-axis, $-1.4$ to $1.4$), with color representing the estimated CAITE magnitude.(a) visualization of RF learner result, (b) visualization of XGB learner result.

### v3

@@ A.10.1 Without mediators
For each scenario , we estimate $\hat{g}_{0}(x)$ and $\hat{g}_{1}(x)$ using randomized clinical trial data. To facilitate visualization, we construct a synthetic grid of covariates where Cov1 and Cov2 vary from 1.5-1.5 to $1.5$ in increments of $0.05$, while all other covariates are drawn from a standard normal distribution. The estimated functions are then applied to this grid to compute the CATTE for each unit. We repeat this process across 100 simulated experiments and compute the average CATTE at each grid point. The resulting surface is interpolated using the R package "akima" and visualized to illustrate the spatial patterns in treatment heterogeneity. We visualize the results under a sample size of $1000$ for each of the four scenarios—Simple, Complex, Global, and Null—using Random Forest and XGBoost as base learners. The visualizations are presented in Appendix Figure A.24 - A.27.
@@
Figure A.24: Visualization of CATTE in the Simple scenario:each panel plots Cov1 (x-axis,1.4-1.4 to $1.4$) and Cov2 (y-axis, 1.4-1.4 to $1.4$), with color representing the estimated CATTE magnitude.(a) visualization of RF learner result, (b) visualization of XGB learner result.
Figure A.25: Visualization of CATTE in the Complex scenario:each panel plots Cov1 (x-axis,1.4-1.4 to $1.4$) and Cov2 (y-axis, 1.4-1.4 to $1.4$), with color representing the estimated CATTE magnitude.(a) visualization of RF learner result, (b) visualization of XGB learner result.
Figure A.26: Visualization of CATTE in the Global scenario:each panel plots Cov1 (x-axis,1.4-1.4 to $1.4$) and Cov2 (y-axis, 1.4-1.4 to $1.4$), with color representing the estimated CATTE magnitude.(a) visualization of RF learner result, (b) visualization of XGB learner result.
Figure A.27: Visualization of CATTE in the Null scenario:each panel plots Cov1 (x-axis,1.4-1.4 to $1.4$) and Cov2 (y-axis, 1.4-1.4 to $1.4$), with color representing the estimated CATTE magnitude.(a) visualization of RF learner result, (b) visualization of XGB learner result.
A.10.2 With a mediator
@@
For each scenario , we estimate $\hat{g}^{Y}_{1}$$\hat{g}^{M}_{1}(x)$ and $\hat{g}^{M}_{0}(x)$ using randomized clinical trial data. To facilitate visualization, we construct a synthetic grid of covariates where Cov1 and Cov2 vary from 1.5-1.5 to $1.5$ in increments of $0.05$, while all other covariates are drawn from a standard normal distribution. The estimated functions are then applied to this grid to compute the CAITE for each unit. We repeat this process across 100 simulated experiments and compute the average CAITE at each grid point. The resulting surface is interpolated using the R package "akima" and visualized to illustrate the spatial patterns in treatment heterogeneity. We visualize the results under a sample size of $1000$ for each of the seven scenarios—Simple-All, Simple-Part, Complex-All, Complex-Part, Simple-Null1, Simple-Null2 and Simple-Global—using Random Forest and XGBoost as base learners. The visualizations are presented in Appendix Figure A.28 - A.34.
@@
Figure A.28: Visualization of CAITE in the Simple-All scenario:each panel plots Cov1 (x-axis,1.4-1.4 to $1.4$) and Cov2 (y-axis, 1.4-1.4 to $1.4$), with color representing the estimated CAITE magnitude.(a) visualization of RF learner result, (b) visualization of XGB learner result.
Figure A.29: Visualization of CAITE in the Simple-Part scenario:each panel plots Cov1 (x-axis,1.4-1.4 to $1.4$) and Cov2 (y-axis, 1.4-1.4 to $1.4$), with color representing the estimated CAITE magnitude.(a) visualization of RF learner result, (b) visualization of XGB learner result.
Figure A.30: Visualization of CAITE in the Complex-All scenario:each panel plots Cov1 (x-axis,1.4-1.4 to $1.4$) and Cov2 (y-axis, 1.4-1.4 to $1.4$), with color representing the estimated CAITE magnitude.(a) visualization of RF learner result, (b) visualization of XGB learner result.
Figure A.31: Visualization of CAITE in the Complex-Part scenario:each panel plots Cov1 (x-axis,1.4-1.4 to $1.4$) and Cov2 (y-axis, 1.4-1.4 to $1.4$), with color representing the estimated CAITE magnitude.(a) visualization of RF learner result, (b) visualization of XGB learner result.
Figure A.32: Visualization of CAITE in the Simple-Null1 scenario:each panel plots Cov1 (x-axis,1.4-1.4 to $1.4$) and Cov2 (y-axis, 1.4-1.4 to $1.4$), with color representing the estimated CAITE magnitude.(a) visualization of RF learner result, (b) visualization of XGB learner result.
Figure A.33: Visualization of CAITE in the Simple-Null2 scenario:each panel plots Cov1 (x-axis,1.4-1.4 to $1.4$) and Cov2 (y-axis, 1.4-1.4 to $1.4$), with color representing the estimated CAITE magnitude.(a) visualization of RF learner result, (b) visualization of XGB learner result.
Figure A.34: Visualization of CAITE in the Simple-Global scenario:each panel plots Cov1 (x-axis,1.4-1.4 to $1.4$) and Cov2 (y-axis, 1.4-1.4 to $1.4$), with color representing the estimated CAITE magnitude.(a) visualization of RF learner result, (b) visualization of XGB learner result.